### v1

# A Personalised Formal Verification Framework for Monitoring Activities of Daily Living of Older Adults Living Independently in Their Homes

Abstract
There is an imperative need to provide quality-of-life to a growing population of older adults living independently. Personalised solutions that focus on the person and take into consideration their preferences and context are key. In this work, we introduce a framework for representing and reasoning about the Activities of Daily Living of older adults living independently at home. The framework integrates data from sensors and contextual information that aggregates semi-structured interviews, home layouts and sociological observations from the participants. We use these data to create formal models, personalised for each participant according to their preferences and context. We formulate requirements that are specific to each individual as properties encoded in Linear Temporal Logic and use a model checker to verify whether each property is satisfied by the model. When a property is violated, a counter example is generated giving the cause of the violation. We demonstrate the framework’s generalisability by applying it to different participants, highlighting its potential to enhance the safety and well-being of older adults ageing in place.
1 Introduction
Advancements in medical science have significantly increased life expectancy in recent decades, resulting in a growing population of older adults (OAs) [1]. As an increasing number of such adults live independently [2], greater attention is being paid to ensuring their safety, addressing their care needs and maintaining their quality-of-life (QoL) [3]. Innovative technologies are increasingly needed to support independent living and enhance the QoL of OAs.
Activities of Daily Living (ADLs) are essential self-care physical tasks that an individual performs on a regular basis. ADLs are fundamental for maintaining health and well-being [4]. Unexpected behaviour during these activities may indicate a decline in cognitive or physical abilities, potentially leading to unsafe conditions [5]. Moreover, monitoring such activities could be crucial for detecting changes that may impact an individual’s QoL. Sensors provide a powerful means of monitoring ADLs and, by extension, detecting deviations, thus potentially facilitating timely interventions that can safeguard or even improve the QoL of OAs.
In this paper, we present a novel framework that combines data from sensors, interviews, home layout and sociological observations (i.e. the systematic analysis of social behaviours and interactions) for representing and reasoning about ADLs in the homes of OAs living independently. The framework leverages unobtrusive, commercially-available sensors to capture ADL events, e.g. opening a refrigerator, closing a shower cubicle door. These events are integrated through an automation platform and stored in a time-series database. To enrich the sensor data, we systematically collect contextual information about daily routines through interviews and sociological observations. This contextual data also enables the customisation of sensor setups and the creation of accurate, person-centred models that account for home layout, privacy preferences, and individual requirements.
This work is specifically aimed at OAs living alone in their homes, with the overall goal of identifying unanticipated behaviours over time.
By modelling the various actions involved in carrying out ADLs as a finite state model [6], we can form an abstraction of the behaviour of the OA. This enables the application of formal methods to verify whether the observed behaviour, captured by a set of sensors, satisfies a set of behavioural properties specific to that OA. These properties are derived from existing general guidelines and from specific requirements of the OA. We use model checking [6], a formal verification technique, to verify whether properties hold for the observed behaviour. More specifically, we use the model checker NuSMV [7] to specify and verify the model, encoding properties in a dialect of Linear Temporal Logic (LTL) [8].
This study builds upon our previous work [9] which involved using public datasets [10]. It significantly extends it through the use of our own data, collected from multiple participants within an OA population in Edinburgh, Scotland. Using the combined sensor and contextual data, we automatically generate models based on the placement of sensors in each participant’s home, while properties are manually encoded to reflect the unique requirements of each participant.
The main contributions of this work are as follows:
1.
A comprehensive framework integrating formal model verification with commercial sensors to capture ADL data in the homes of OAs. This is then extended with explicit contextual data. This facilitates the creation of personalised models and property specifications, linking ADL sensor events to OA preferences and needs.
2.
A rigorous and systematic approach for the specification of models and properties based on information acquired from semi-structured interviews with each participant
3.
Our approach is generalisable and we have applied it to a diverse range of OAs. Here we illustrate with two participants with contrasting ADLs.
This paper is organised as follows: Section 2 discusses related work in sensor-based ADL monitoring and formal methods. Section 3 describes the main components of the framework and data collection. Section 4 describes the data processing and personalised models. Section 5 describes the experiments and the results. In Section 6 we present our conclusions and outline future work.
2 Related Work
Early work on ADL monitoring using unobtrusive sensors has primarily focused on recognising and tracking ADLs. Initial studies, such as those by Munguia et al. [11, 12], used contact sensors and energy switches to capture events involving people living alone to track daily activities. Following this, a range of studies have examined various aspects of ADL monitoring. For instance, Kang et al. [13] focused on the use of IoT devices and sensor fusion to capture ADL events and improve data reliability, while Matsui et al. [14] proposed semi-automated mechanisms that relied on physical buttons for activity annotation. Tonkin et al. [15] worked on a multi-sensor dataset collected in a controlled residential setting, enhancing its accessibility through improved documentation, annotations, and organisation. That dataset comprises short-duration scripted ADLs. The use of formal methods in ADL monitoring has also been explored in previous studies. Konios et al. [16] used Petri nets to distinguish between normal and abnormal behaviour in data captured from motion, light, temperature, and contact sensors. Garcia-Constantino et al. [17, 18] collected data from ambient and wearable sensors to identify abnormal behaviour during preparation of beverages in the kitchen using Petri nets. Magherini et al. [19] used temporal logic to automate the recognition of three ADLs within a hypothetical scenario involving a person living alone. The activities assumed specific sequences and timing of actions, which were captured by camera sensors.
A shared characteristic of the above approaches is their reliance on common assumptions about the ADLs a person performs and where these activities are conducted. Moreover, the aforementioned studies, involving formal verification and ADLs, were carried out in controlled, simulated environments. In contrast, our work focuses on OAs performing ADLs in their own homes and personalises the approach so that the data collection, models and properties to be verified are all tailored to their situation and preferences, an approach that has been advocated by other studies [20, 21]. Furthermore, we streamline our pipeline by using an automated tool that generates formal symbolic models based on home layouts and the positioning of sensors.
3 Framework Components and Data Collection
This section provides an overview of the main components used in our framework and the data collection process.
3.1 Framework Overview
The main aspects of our framework are illustrated in Figure 1. They are categorised into two groups of elements: 1) those involved during data collection in the participant’s home and 2) those used after data collection. The components present during data collection are:
Participants: Our study is a subject-centric approach. Through semi-structured interviews, participants provide insights into their daily routines, preferences and specific needs, which will later be expressed as LTL properties. This also allows us to personalise the placement of sensors, by considering individual needs, their living environment and other aspects that may otherwise be overlooked (see Section 3.2).
Sensors: Two types of sensors are deployed in each participant’s home: contact and motion. Contact sensors monitor states related to doors, windows and appliances, providing data on object interactions. Motion sensors detect movement or changes in the position of participants within specific areas of the home. All sensors operate over the Zigbee wireless communication protocol [22].
Sensor Platform and Storage: We employ open-source solutions for sensor data acquisition and management. Home Assistant [23], a home automation platform, orchestrates data acquisition from our sensors. This provides a flexible and extensible setup for data aggregation. We store the sensor-generated data in InfluxDB [24], a high-performance time-series database optimised for handling large volumes of time-stamped data.
The components utilised post data collection are:
Interview Analysis: Interviews are parsed to identify participants’ routines and preferences. Qualitative analysis informs property specification, highlighting relevant aspects of personalisation, such as environment and needs (see Section 4.1).
Model Generator: This generates a formal model of the participant’s home environment. It incorporates a representation of their home layout, sensor placement within the layout and the observed behaviour captured by the sensors (see Section 4.2).
Properties: For each participant we define a set of properties which will be formally checked. These properties are derived from existing safety guidelines and the individual’s contextual information, including needs and preferences , creating requirements for the expected behaviour. We encode the properties in LTL and use a model checker to verify that the model satisfies the specified requirements (see Section 4.3).
NuSMV: This is the model-checking engine. It verifies the correctness of the formal model against specified properties and generates a sensor event trace (cf. a counterexample) highlighting the sequence of events leading to any violation.
3.2 Data Collection Process
Each potential participant is provided with an information sheet outlining the purpose and scope of the study. Consent forms are used to ensure participants formally agree to taking part and that they understand their rights to withdraw at a later date if they so desire. Ethical approval for the study protocol was obtained from the School of Informatics Ethics Committee (application number 558534) at the University of Edinburgh prior to the start of recruitment.
In what follows, we describe the qualitative and sensor data collection. For different participants, data are collected over periods of one to three weeks depending on their availability.
3.2.1 Qualitative Data
Once consent is obtained, a home visit is scheduled with the participant for the collection of qualitative data. A semi-structured interview is used by the research team to gauge the participant’s daily routines, the context in which these are performed and the needs and preferences the participant may have prior to sensors installation. The interview guide is divided into five sections: i) ADLs, ii) context and environment including the impact of sensors on the daily routine, iii) thoughts on privacy and security, iv) current needs and preferences, feedback and suggestions and v) general well-being. From the interview and observations, we also obtain insight into the possible placement of sensors. For example, some participants express discomfort with motion sensors in private areas such as bedrooms, but are comfortable with contact sensors in the same room. Figure 2a shows an example of the layout and the placement of sensors for one participant, based on their activities and privacy concerns. At the end of sensor data collection, a second interview is conducted to assess any changes in the participant’s routines or needs and to determine whether their behaviour was influenced by the presence of sensors. Qualitative data are used to identify requirements that can be translated into properties regarding the participant’s behaviour.
3.2.2 Sensor Data
Sensor data are collected from contact and motion sensors placed throughout the participant’s home. The latter are positioned to cover living spaces and areas of transit, while contact sensors are attached to objects like cupboard doors or drawers. Over the course of our study, the number of sensors installed in homes ranged from to , while data collection ranged from to days. Passive infrared (PIR) sensors are used detect movement, with a detection angle of and a range of up to metres. These sensors have two possible states: clear and detected. If motion is registered while the sensor is in a clear state, the state changes to detected and remains in that state for seconds. If no further motion is registered within seconds, the sensor switches back to the clear state. If motion is registered while the sensor is in the detected state, the 30-second timer is restarted. Each contact sensor has two possible states: open and closed. The contact sensor consists of two components: body and magnet. When the body is within mm of the magnet, the state of the sensor is closed and beyond this distance, the state is open. Additionally, each contact sensor registers its internal temperature. Figure 3a shows a motion sensor used to detect movement in the area where medication is kept on a shelf in a participant’s kitchen. Figure 3b shows contact sensors attached to various drawers in the participant’s kitchen. Both motion and contact sensors operate on an event-driven basis and are connected to a ConBee II Zigbee gateway. As previously mentioned, we use Home Assistant to orchestrate data acquisition from the sensors and store our data using InfluxDB. The automation platform, gateway and database run on an NVIDIA Jetson Nano with a quad-core ARM CPU, a -core GPU, and GB of RAM, operating on a Linux-based system. The collected data are exported from the database in CSV format, resulting in dataset sizes ranging from MB to MB.
4 Data Processing and Models
In this section we describe the data processing steps, the encoding of properties and the generation of the symbolic models.
4.1 Preprocessing
4.1.1 Information from participants
Each participant is first allocated a unique identifier to ensure privacy and confidentiality. The mapping between participants’ personal details and their unique identifiers is kept separately under password protection and with controlled access.
The data generated from the interviews allow us to learn about each individual’s ADLs, personal preferences and living context, which, in turn, help us model the properties that reflect general and specific ADLs. We apply Framework Analysis [25], a qualitative analysis method, to reveal patterns and themes in the interview data using NVivo software (Limivero, v14) [26]. We identify properties of ADLs by considering both our questions during interviews and the routines participants highlighted as important. These properties are then categorised as general (common to everyone, like eating or washing), individual-specific (like pet care or hobbies), or routine but non-daily activities that impact health (like using medication). This approach allows us to have a detailed representation of the properties used to verify general and specific ADLs.
4.1.2 Sensors
Data from contact and motion sensors are first processed to reduce noise and retain events that reflect user interactions.
A contact sensor broadcasts its current state every time it changes from open to closed, closed to open or when there is a fluctuation of in the device’s temperature. Repetitions of the same state are taken to be noise in the data triggered by temperature fluctuations so we remove consecutive open (resp. closed) states and consider only the first occurrence. This provides a more accurate representation of the participant’s interaction.
A motion sensor broadcasts its state whenever there is a transition between clear and detected states. If the state is detected and no additional motion is registered within seconds, the sensor broadcasts an artificial clear state after the seconds. We remove clear states from the data as they do not reflect the actual cessation of motion. This provides a more accurate representation of movement, treating motion sensor events as instantaneous. In addition, if continuous clear and detected states occur for a motion sensor, with no changes in other sensors, we only consider the initial detected state of the motion sensor and discard the subsequent states.
4.2 Model generator
We use NuSMV to specify a model that represents the home and behaviour of the participant. The model is a transition system where changes in sensors can occur one at a time. This is in line with what we observe in the data. The model consists of: i) states, which represent possible configurations of sensor states and where the participant is during the execution of the system; ii) transitions, that correspond to events, which describe how the system evolves from one state to another; and iii) the initial conditions that define the sensors’ starting configuration. The model is verified against a set of logically-encoded properties (see Section 4.3).
When a property is violated, NuSMV generates a counter-example: a specific sequence of states that shows the conditions leading to the violation, making it easier to identify and understand its cause.
To produce a bespoke model for a participant we use our model generator, which takes as input a file that specifies the sensors installed and the layout of a participant’s home. For each sensor, we specify its type, installation room, specific location and the name of the sensor in the model. Table 1 presents a sample of the sensors deployed in the home previously depicted in Figure 2a. The input file also includes the abstract layout of the home that describes the spatial arrangement of the rooms. This corresponds to a directed graph, with each node representing a room and the edges representing the connections between rooms. Figure 2b depicts the graph of the layout in Figure 2a.
| Type | Room | Location | Name |
| Contact | Kitchen | Rubbish bin lid | RubbishLid |
| Contact | Kitchen | Refrigerator door | RefrigeratorDoor |
| Contact | Kitchen | Microwave door | MicrowaveDoor |
| Motion | Kitchen | Facing cabinets | MotionKitchen |
| Contact | Kitchen | Drawer cutlery | DrawerCutlery |
| Contact | Kitchen | Cupboard food | CupboardFood |
| Motion | Corridor | Facing corridor | MotionCorridor |
| Contact | Bathroom | Shower cubicle door | ShowerDoor |
| Motion | Bedroom | Facing bed | MotionBedroom |
The behaviour of each sensor in the model is encapsulated in a module. An excerpt of the NuSMV code for the contact module is shown in Listing 1. For each contact sensor we define three variables: its state, activation and deactivation, which are initialised as FALSE (see lines - ). In the module, the state of a contact sensor is a boolean variable where FALSE indicates the body and magnet of the sensor are nearby (e.g. drawers closed as shown in Figure 3b) and TRUE indicates they are apart. The activated boolean variable is set to TRUE at the time step when the state transitions from FALSE ( state) to TRUE (state), as shown in line , and it remains TRUE for one time step only. Note that while activated represents interaction with the person and thus serves as evidence of their activities, the state of a sensor being open – which can last for more than one time step – does not. The deactivated boolean variable is set to TRUE at the time step when the state transitions in the opposite direction, as shown in line , also remaining TRUE for one time step only. Similarly, while deactivated represents interaction by a participant, the state of a sensor being closed does not.
The behaviour of a motion sensor in the model follows a similar structure but with two variables instead of three: its state and activation. The active boolean variable is set to TRUE at the time step when the state transitions from FALSE to TRUE and it remains TRUE for one time step only. However, unlike the contact sensor, the state is explicitly set to FALSE immediately after being TRUE.
We provide the model with a subtrace of events that were actually captured by the different sensors. This subtrace represents the observed behaviour of a participant during a specific period of time. Consider a trace for a set of sensors . Given a pair of unique time markers, and – manually inserted to mark the start and end of a time interval – the subtrace is the set of all events, chronologically ordered, from to , where denotes the time of an event:
In the model, an -dimensional vector represents the state of all sensors in a home at time . For example, Figure 4 shows a subtrace in which a total of three sensors , and change states over time. In this case, at time , sensors and are deactivated while sensor is active in the corresponding 3-dimensional vector .
Using subtraces and the vector of states enables the analysis of the observed behaviour over specific intervals, helping mitigate the state space explosion problem, which refers to the rapid growth of possible system states in NuSMV. These intervals enable the identification of the specific times at which the behaviour was verified.
The model specifies the configuration of a group of sensors (such as in a room) by using a subvector of . This allows us to capture a change in the room/group as a difference in the subvector in subsequent time steps. For example, grouping sensors including the shower cubicle door and the bathroom motion under the name Bathroom enables us to express a change in all these sensors as a single variable, , in our model. A detailed explanation of the various parts of a NuSMV model can be found in our previous work [9].
4.3 Formalism and Property Specification
We are interested in modelling behaviour over time. This makes the temporal logic LTL apt for our work as it allows us to specify properties using various temporal operators. In our work, we use LTL extended with past-time operators [27], henceforward referred as past Linear Temporal Logic (pLTL), as it enables reasoning about events that have already occurred. This extension offers practical expressiveness, allowing properties based on sequencing of events to be specified more clearly and compactly.
We construct pLTL formulas, representing properties for a given participant (see next section), and then check whether all execution paths (sequences of states) of the model satisfy them. A formula consists of atomic propositions (a property inherent in the state of the model), propositional operators and temporal operators. Formally, formulas are generated according to the following BNF grammar:
Satisfaction of a formula by a path at time index is denoted by . An atomic proposition is satisfied if the state at that index of the trace is labelled with it. The non-temporal logical operators , , and have the same meaning as in propositional logic.
Some of the temporal operators of pLTL are described in Table 2. This selection does not preclude the use of other operators; rather, it reflects those that are specifically chosen for the properties discussed here, recognising that different properties may require different operators.
4.4 Behaviour-related Properties for Participants
Properties are identified and categorised using the analysis of qualitative data mentioned in Section 4.1. In what follows, we present three of these in case studies pertaining to two participants, referred to as participant A and participant B. Each participant is in their late 60s, retired, living independently, not receiving external care and follows defined daily routines. This selection of properties does not preclude the inclusion of others; however, space constraints limit what we can present here.
| Operator | Interpretation |
| Next () | In , is expected to hold at the immediate next time step |
| Previous () | In , held at the time step immediately before the current one |
| Eventually () | In , will hold at some point in the future, though not necessarily immediately |
| Once () | In , held at some point in the past, at or before the current time step |
| Globally () | In , holds at the current time and continues to hold at all future time steps |
| Until () | In , holds continuously up to a future point when becomes true. At that point must hold and must have held at every earlier step |
A daily routine corresponds to a subtrace delimited by time markers startActivity and endActivity. To improve clarity and reduce repetition in the specifications, we define the following pattern:
This expression specifies that, once the routine begins (), but before the routine ends (), the condition must eventually hold.
The following two properties describe Participant A’s early morning routine, which takes place in the home depicted in Figure 2.
Property 1
This specifies that participant A moves from the bedroom to the bathroom to take a shower. Since their preference is to do this after waking up, we check this property in accordance to their morning behaviour (see Section 5). The property falls under the general category of the qualitative analysis. For the participant, performing this ADL involves moving from the bedroom to the corridor, then to the bathroom and closing the shower cubicle door. The property (1) specifies that motion in the bedroom () holds until there is motion in the corridor (), which in turn holds until there is a change in the bathroom (). The change in the bathroom must persist until the shower cubicle door is closed (). The bedroom and the corridor each have a single motion sensor; in both cases, one sensor was sufficient to cover the entire area, given the layout. The bathroom has a motion sensor and a contact sensor. A change in the bathroom can be triggered by motion detected in the bathroom () or by closing () or opening () the shower cubicle door.
Note that the property allows the participant to open the shower cubicle door (if closed), enter the cubicle, and then close the door to take a shower. This accounts for situations where the door may have been closed earlier, for example, after cleaning the shower.
Property 2
This specifies that participant A opens the refrigerator to get the salad to feed their pets (before interacting with any other cupboards). Note that since their preference is to do this before breakfast, we check this property in accordance to their morning behaviour (see Section 5). From the qualitative analysis this is an individual-specific property. The property (2) specifies that if the refrigerator door is open () and the door has not been opened in any previous steps (), then none of the drawers or cupboards have been previously opened or closed (), where KitchenCupboards represents all drawers and cupboards in the kitchen.
Note that this property allows the participant to move in the kitchen or to open or close the rubbish bin or the kitchen door, each having a contact sensor, without violating the property.
Property 3
This corresponds to the routine category of the qualitative analysis. It is formulated for participant B and describes the participant’s interaction with the medication dispenser located in the kitchen each morning within a defined time window specified by the participant, as detailed in the experiments (see Section 5). This dispenser is monitored by a motion sensor, as shown in Figure 3a. To reach the kitchen, the participant must first pass through the lounge. The participant may move between the lounge and kitchen, and can interact with various sensor-equipped fittings before reaching the dispenser. The property (3) specifies that change in the lounge () holds until there is change in the kitchen () and that change in the lounge or kitchen () holds until the dispenser is accessed ().
Additional Properties
We also encoded properties specifying the allowable sequence of events imposed by physical constraints. This was done to ensure the sensor events align with the physical layout, thereby verifying correct sensor functioning. We focus here on the physical layout for Participant A’s home (see Figure 2). Property (4) specifies that if motion is detected in the bedroom and, at a later time, in the bathroom, motion must also be detected in the corridor in between.
Property (5) specifies that when the participant moves from the lounge to the kitchen, the first detected event in the kitchen must be motion.
5 Experiments, Results and Analysis
In this section we present the results of the experiments conducted to verify whether the properties described in Section 4.3 are satisfied for the observed behaviour of the participants. Each property is individually evaluated on a model generated for the participant. The model includes the data captured from sensors (i.e. the subtrace) during the period of time identified as relevant for the property (e.g. the morning for Property 1).
We use a sliding time window to segment the data into minute subtraces (see Section 4.2), which are then passed to the model, allowing for more precise identification of when a property holds or fails. To avoid cutting off a subtrace where a property might hold, we allow a minute overlap between subtraces. The time window was chosen pragmatically, as the identified ADLs were of short duration. Using shorter time windows also reduces the number of events per subtrace, helping to mitigate the known challenge of state space explosion problem and improve performance.
From the qualitative analysis, we identify that participant A performs their morning activities around AM every morning. To compensate for early and late mornings, we consider all data captured between AM and AM when evaluating the properties. Similarly, participant B performs their activities around AM every morning; we consider all data captured between AM and AM.
Table 3 summarises the outputs when verifying Property 1 and Property 2 for participant A. The table is organised with grouped columns for each property. The Status column states whether the property was satisfied or violated within the morning routine. The Time column indicates the time frame for which the property is satisfied. The Seq column shows the number of subtraces identified within the interval frame for which the property is satisfied.
Property 1 was satisfied on all dates between AM and AM except August . Analysis of the counterexample revealed that participant A followed the expected sequence of events between AM and AM, but once in the bathroom, returned to the corridor before going back to the bathroom and closing the shower cubicle door. This unexpected transition caused the property to be evaluated as FALSE, even though it is likely that the activity of taking a shower in the morning was completed. This suggests that temporal sequences may need refinement to accommodate minor deviations in daily routines.
Similarly, Property 2 was violated only on Sunday, August, when participant A did not open the refrigerator before AM. On the same day, a slight delay in the shower routine was observed, indicating that weekend routines could shift in terms of when they occur. This suggests that defining the start of the day relative to the participant’s first set of sensor activations is an aspect worth exploring further. In addition a higher level of specificity may be needed when conducting interviews to address such nuances.
| Date | Property (morning shower) | Property (refrigerator) | ||||
| Status | Time | Seq | Status | Time | Seq | |
| 23-08-24 | Satisf. | 7:50–8:20 | 2 | Satisf. | 7:50–8:20 | 2 |
| 24-08-24 | Satisf. | 8:30–9:00 | 2 | Satisf. | 9:00–9:30 | 2 |
| 25-08-24 | Satisf. | 8:50–9:10 | 1 | Viol. | – | 0 |
| 26-08-24 | Satisf. | 8:20–8:50 | 2 | Satisf. | 7:30–9:30 | 1 |
| 27-08-24 | Satisf. | 8:10–8:40 | 2 | Satisf. | 7:30–9:30 | 2 |
| 28-08-24 | Viol. | – | 0 | Satisf. | 7:30–9:30 | 2 |
| 29-08-24 | Satisf. | 8:30–8:50 | 1 | Satisf. | 7:30–9:30 | 2 |
The two physical properties specified for the layout of participant A were evaluated throughout the day from AM to PM. Property 4 held throughout the day. Property 5, related to motion in the kitchen, was occasionally violated. The violations occurred when the participant entered the kitchen and opened the bin. We believe that these occasional violations occurred because motion near the bin was not properly captured by the motion sensor in the kitchen.
For participant B, we evaluated the property that the medication dispenser was accessed while the participant moved between the kitchen and the lounge between and . Table 4 shows the dates and time intervals for which the property was satisfied. The property was not satisfied on Friday November ; participant B entered the lounge, a necessary step to reach the medication, only after AM. While some ADLs allow for flexibility in when they occur, the intake of medication is a more rigid activity. A deviation in this activity potentially implies a health safety risk.
| Date | Status | Time | Seq |
| 16-11-24 | Satisf. | 7:50 - 8:50 | 4 |
| 17-11-24 | Satisf. | 8:20 - 8:50 | 2 |
| 18-11-24 | Satisf. | 8:40 - 9:00 | 1 |
| 19-11-24 | Satisf. | 7:20 - 8:20 | 5 |
| 20-11-24 | Satisf. | 8:10 - 9:00 | 4 |
| 21-11-24 | Satisf. | 7:00 - 7:40 | 3 |
| 22-11-24 | Viol. | – | 0 |
6 Conclusion
In this paper, we have presented a framework for representing and reasoning about ADLs performed by older adults. As part of our work, we recruited participants aged to . For each participant, we collected both qualitative and quantitative data over a period ranging from one to three weeks. Our approach leverages contextual information from the participants and data captured by commercially available sensors installed in their homes to build tailored models. We use the qualitative data to specify properties that are tailored to the participants’ preferences and needs. Properties are encoded using LTL extended with past-time operators. Models are automatically generated for each participant and the model checker NuSMV is used to evaluate whether the observed behaviour of participants conform to the encoded properties. To demonstrate the applicability of our framework, we conducted experiments on real-world datasets from two participant living independently. The results support the feasibility of our framework for analysing tailored properties and identifying deviations from expected behaviour. From the case studies presented, we observe that one deviation corresponds to changes in a participant’s routine when taking a shower. Another deviation was triggered because the activity (medication intake) was not performed. Further analysis reveals that, although the activities are not performed at the expected time, they are carried out correctly, albeit a few minutes later. Although this may suggest that the timing for property verification could be relaxed, timely performance of activities such as medication intake might be essential and any violation indicate the need for an intervention.
Acknowledgements
This research was funded by the Legal & General Group (research grant to establish the independent Advanced Care Research Centre at the University of Edinburgh). The funder had no role in conducting the study, interpretation or the decision to submit for publication. The views expressed are those of the authors and not necessarily those of Legal & General.